We define a function by:

f(x) = { x +1 si -1 < x < 0; si 0 < x <1, 1
and we suppose that f (x + 2) = f (x) for all x ∈ R.
(a) Draw the graph of the function f (x) over the interval [−3, 3].
(b) Find the Fourier series for the function f (x).
Answer #1

X+1 fle) = -1 $220 1 osx1 f(x+2) = f(x) VXEIR period of f(x) = 2 2 2 2 =) La 1 1 2 3 b where Op T זמ graph of hurlion f(x) ovbn 2 K S F(X) Sin ( mTlx) dx -L s f(x) yin 6 2 nax) dx / $(x+1) sin (nux) + Si.siu (noix de -1/17 + (1-(-10%) = -(-10% EMA f(